# Foundation Models for Anomaly Detection: Vision and Challenges

### v1

@@ 4.1 Serialization-based Detection
Myriad Li et al. (2023) GPT-3.5 LVLM Image Industry Link
VAD-LLaMA Lv and Sun (2024) LLaMA LVLM Video Real-world surveillance -
LogConfigLocalizer Shan et al. (2024) GPT-4 LLM Log Software system Link
Others LogiCode Zhang et al. (2024b) GPT-4 LLM Image Industry -
AnomalyRuler Yang et al. (2025) GPT-4 & Mistral-7B LVLM&LLM Video Real-world surveillance Link
Audit-LLM Song et al. (2024) GPT-3.5 LLM Log Audits -
@@ 4.2 Encoding-based Detection
Deep learning has demonstrated impressive capabilities in understanding different data structures through neural networks, which excel at recognizing patterns in complex and high-dimensional data. As illustrated in Figure 3(b), encoding-based detection leverages the advantages of deep learning to incorporate complex characteristics present in source data, allowing FMs to be characteristic-aware:
@@
where $f_{DL}$ is the encoder for data representation learning based on deep learning (DL). DL-based prediction also utilizes a parser to retrieve FM output.
To combat the challenge that traditional industrial anomaly detection models can only provide anomaly scores of data samples and the threshold must be manually determined, AnomalyGPT Gu et al. (2024) employs an image decoder to provide fine-grained semantics and design a prompt learner to fine-tune the LVLM using prompt embeddings. Therefore, it could directly assess the locations of anomalies and provide image information in a human-understandable manner.
LAVAD Zanella et al. (2024) uses LLMs to detect video anomalies exclusively from a scene description without further training. After generating a textual description for each video frame with a pre-trained VLM, an LLM is applied to capture the dynamics of the scene and summarize captions within a temporal window, which will be further used to provide an anomaly score for each frame.
Discussion. Utilizing FMs directly as detectors shows superiority in processing textual attributes of different types of data, especially achieving remarkable zero-shot performance compared with traditional DL models. The ultimate goal is to develop and refine methods for encoding these structured information into a format that FMs can comprehend and manipulate effectively and efficiently. Despite the fact that LLMs have shown their potential in understanding long-context and mathematical reasoning, they cannot be directly applied as an anomaly detector without prompt learning and fine-tuning strategy. Moreover, for encoding-based detection, training an additional DL module and inserting it into FMs for joint training is challenging due to the problem of vanishing gradients in the early layers of deep transformers Li et al. (2024b).
@@ 5.1 Detection-based Explanation
Direct detection. Holmes-VAD Zhang et al. (2024a) is composed of three key components, namely, a Video Encoder for encoding input video, a Temporal Sampler for abnormal frame prediction, and a Multi-modal LLM for generating text explanations. To create a generic model capable of detecting semantic anomalies in business processes, DABL Guan et al. (2024) fine-tunes the Llama by incorporating traces into question and answer content. Specifically, every trace item (i.e., business process) and its label are inputs to the foundation model so that the model could be fine-tuned in a supervised way. Then, the fine-tuned model outputs whether the given trace is normal or anomalous, and provides causes of anomalous traces. LLMAD Liu et al. (2024b) is an LLM-based framework for time series anomaly detection. It improves detection performance and interpretation quality by injecting data background and domain knowledge into model via time-series In-Context Learning and Chain-of-Thought approach guiding its decision-making process.
Indirect detection. In Myriad Li et al. (2023), a vision expert tokenizer embeds the anomaly map into vision expert tokens to make LLM perceive prior knowledge, and a vision expert instructor generates domain vision-language tokens to compensate for the errors of vision experts. VAD-LLaMA Lv and Sun (2024) is composed of a video encoder and a foundation model, which are connected by a projection layer called adapter. By treating the foundation model as a question-answer system, the foundation model provides detailed information (e.g., the concrete video content and the accurate time frames when anomalies appeared) of anomalies by answering the human queries.
@@ 5.2 Verification-based Explanation
To further validate the authenticity of detected anomalies,  Park (2024) proposes an LLM-based multi-agent framework to analyze the detailed anomaly information, which serves as a critical interface between the AI-driven analysis process and human decision-making. With the aim of assisting end-users in coping with configuration errors in software system through log analysis,  Shan et al. (2024) proposes a two-stage strategy to localize the root-cause configuration logs. The first stage is to identify anomalous logs through obtaining key log messages, and the second stage is to use an LLM for further verification with the help of their strong power in natural language understanding and processing. Specifically, strategies are designed to increase the reliability of LLMs’ judgments and provide additional information about the configuration errors.
Discussions. Although FMs excel in providing explanations on the anomaly detection results, its interpretability is still limited due to the difficulty of well-crafted prompts. Actually, creating these prompts needs domain specific knowledge from experts. Therefore, how to improve the data quality with acceptable labeling costs is still a core challenge to be solved in the future.
@@ 6 Open Challenges and Way Forward
Based on the above review and analysis, we believe that there are still many challenges and potentials for further enhancement in this field. In this section, we list some future research directions for further exploration.
@@ 6.1 Efficiency
Notwithstanding the stellar progress and accomplishments of FMs, the improvements of performance in these models come at the expense of the model’s efficiency. However, anomaly detection models need to be particularly efficient because they are often applied in real-time, high-stakes environments where timely detection and response to anomalies are critical. Common application scenarios include fraud detection and healthcare, where even small delays of identification may lead to significant financial loss or even safety risks. While some lightweight adaptation techniques have been proposed to improve the parameter efficiency, Lester et al. (2021) gives an example that, as the model size increases, the performance gap between full fine-tuning and lightweight adaptation is diminishing rapidly. Therefore, exploring mechanisms to balance the trade-off between efficiency and expressivity of FMs still remains a notable challenge.
@@ 6.2 Bias
In recent years, FMs have led to an extraordinary level of homogenization, which refers to the consolidation of methodologies for building machine learning systems across a wide range of applications Bommasani et al. (2021): current NLP models are often adapted from one of popular FMs like BERT. Despite that any improvements of FMs can help produce immediate benefits, such kind of extension might also have the potential to inherit or even amplify the problematic biases of these models. A biased anomaly detection model may potentially lead to incorrect or unfair outcomes. For example, a biased model may disproportionately flag specific data points or groups (e.g., gender and racial bias) as anomalous even when they are normal, limiting their capability of detecting genuine anomalies. Such kind of false positives or false negatives can lead to unfair treatment of certain individuals or groups, missed critical alerts, or a lack of trust in the system, ultimately compromising its effectiveness and fairness. Therefore, it is critical to eliminate the intrinsic bias present within FMs, thereby further avoiding extrinsic harms in downstream anomaly detection systems.
@@ 6.3 Explainability
Despite that FMs are currently used as interpreters, in many works, to illustrate why anomalies are detected by providing textual explanations in a human-understandable manner, the models themselves are not transparent and interpretable enough to give explanations of the internal model structures and behaviors. However, providing evidence and logical steps for decision-making is a critical issue in anomaly detection especially in applications like healthcare and finance, where understanding the reasoning behind an anomaly is as important as detecting it Chandola et al. (2009). Therefore, how to improve the interpretability of FMs remains an open research question, and it is absolutely important for researchers to enhance the trust, usability, and decision-making of anomaly detection.
@@ 6.4 Multimodality
While multimodality is considered a critical element of intelligence, and serves as a crucial component for the development of both thorough and broad comprehension of the world, models that go beyond simple alignment of vision and language are yet to emerge Radford et al. (2021); Lu et al. (2019). In real-world application scenarios, taking healthcare as an example, medical data are highly multimodal, with various data types, scales, and styles. Multimodal models generally have better performance on anomaly detection by integrating and analyzing data from different sources than models with a single modality. However, current anomaly detection models are mainly developed for single modality (e.g., text, image, and gene), and do not learn from various modalities. By harnessing information across different data types, multimodal data can provide richer and more robust anomaly detection, thereby improving the accuracy of detection results across various applications. Future studies should further examine the design of FMs that integrate various modalities and domains.
@@ 7 Conclusion
The application of FMs to anomaly detection has emerged as a prominent area of research in recent years. In this survey, we provided an in-depth review of the use of FMs in anomaly detection tasks. We introduced a novel taxonomy classifying the methods into three approaches based on the role of FMs played in different stages of anomaly detection, namely encoder, detector, and interpreter. This taxonomy clarifies how FMs can empower the process of data representations, anomaly detection, and explainable result analysis of detection. We also discussed challenges and highlighted several future research directions, aiming to shed light on the advances and challenges in the field of anomaly detection with FMs, thereby encouraging further progress in this domain. As FMs continue to evolve, their role in anomaly detection is expected to expand, unlocking new possibilities for more accurate, interpretable, and scalable anomaly detection systems.

### v2

@@ 4.1 Serialization-based Detection
Myriad (li2023myriad, ) GPT-3.5 LVLM Image Industry Link
VAD-LLaMA (lv2024video, ) LLaMA LVLM Video Real-world surveillance -
LogConfigLocalizer (shan2024face, ) GPT-4 LLM Log Software system Link
Others LogiCode (zhang2024logicode, ) GPT-4 LLM Image Industry -
AnomalyRuler (yang2025follow, ) GPT-4 & Mistral-7B LVLM&LLM Video Real-world surveillance Link
Audit-LLM (song2024audit, ) GPT-3.5 LLM Log Audits -
@@ 4.2 Encoding-based Detection
Deep learning (DL) has demonstrated impressive capabilities in understanding different data structures through neural networks, which excel at recognizing patterns in complex and high-dimensional data. As illustrated in Figure 3(b), encoding-based detection leverages the advantages of deep learning to incorporate complex characteristics present in source data, allowing FMs to be characteristic-aware:
@@
where $f_{DL}$ is the encoder for data representation learning based on deep learning. DL-based prediction also utilizes a parser to retrieve FM output.
To combat the challenge that traditional industrial anomaly detection models can only provide anomaly scores of data samples and the threshold must be manually determined, AnomalyGPT (gu2024anomalygpt, ) employs an image decoder to provide fine-grained semantics and introduces a prompt learner to fine-tune the LVLM using prompt embeddings. Therefore, it can directly assess the locations of anomalies and provide image information in a human-understandable manner.
LAVAD (zanella2024harnessing, ) uses LLMs to detect video anomalies exclusively from a scene description without further training. After generating a textual description for each video frame with a pre-trained VLM, an LLM is applied to capture the dynamics of the scene and summarize captions within a temporal window, which will be further used to provide an anomaly score for each frame.
Discussion. Utilizing FMs directly as detectors shows superiority in processing textual attributes of different types of data, especially achieving remarkable zero/few-shot performance compared with traditional DL models. The ultimate goal is to develop and refine methods for encoding these structured information into a format that FMs can comprehend and manipulate effectively and efficiently. Despite the fact that LLMs have shown their potential in understanding long-context and mathematical reasoning, they cannot be directly applied as an anomaly detector without prompt learning and fine-tuning strategy. Moreover, FMs tend to miss or normalize anomalies, especially some subtle or domain-specific ones, because they are trained on large corpora of normal data without specific training stages (russell2024aad, ). As for encoding-based detection, training an additional DL module and inserting it into FMs for joint training is challenging due to the problem of vanishing gradients in the early layers of deep transformers (li2024survey, ).
@@ 5.1 Detection-based Explanation
Direct detection. Holmes-VAD (zhang2024holmes, ) is composed of three key components, namely, a Video Encoder for encoding input video, a Temporal Sampler for abnormal frame prediction, and a Multi-modal LLM for generating text explanations. To create a generic model capable of detecting semantic anomalies in business processes, DABL (guan2025dabl, ) fine-tunes the Llama by incorporating traces into question and answer content. Specifically, every trace item (i.e., business process) and its label are inputs to the foundation model so that the model could be fine-tuned in a supervised way. Then, the fine-tuned model outputs whether the given trace is normal or anomalous, and provides causes of anomalous traces. LLMAD (liu2024large2, ) is an LLM-based framework for time series anomaly detection. It improves detection performance and interpretation quality by injecting data background and domain knowledge into model via time-series In-Context Learning and Chain-of-Thought approach guiding its decision-making process.
Indirect detection. In Myriad (li2023myriad, ), a vision expert tokenizer embeds the anomaly map into vision expert tokens to make LLM perceive prior knowledge, and a vision expert instructor generates domain vision-language tokens to compensate for the errors of vision experts. VAD-LLaMA (lv2024video, ) is composed of a video encoder and a foundation model, which are connected by a projection layer called adapter. By treating the foundation model as a question-answer system, the foundation model provides detailed information (e.g., the concrete video content and the accurate time frames when anomalies appeared) of anomalies by answering the human queries.
@@ 5.2 Verification-based Explanation
To further validate the authenticity of detected anomalies,  (park2024enhancing, ) proposes an LLM-based multi-agent framework to analyze the detailed anomaly information, which serves as a critical interface between the AI-driven analysis process and human decision-making. With the aim of assisting end-users in coping with configuration errors in software systems through log analysis,  (shan2024face, ) proposes a two-stage strategy to localize the root-cause configuration logs. The first stage is to identify anomalous logs through obtaining key log messages, and the second stage is to use an LLM for further verification with the help of their strong power in natural language understanding and processing. Specifically, strategies are designed to increase the reliability of LLMs’ judgments and provide additional information about the configuration errors.
Discussions. It is acknowledged that providing human-understanable explanations for anomaly detection results is useful for high-stake domains like healthcare and finance, where domain experts without AI knowledge can engage with the system. Apart from the explanation results, incorporating external knowledge that are not encoded in the raw data has the potential to reduce false positives through context (zhang2024logicode, ).
Although FMs excel in providing explanations on the anomaly detection results, its interpretability is still limited due to the difficulty of well-crafted prompts. Actually, creating these prompts needs domain specific knowledge from experts (cao2024domain, ). Therefore, how to improve the data quality with acceptable labeling costs is still a core challenge to be solved in the future.
@@ 6 Open Challenges and Way Forward
Based on the above review and analysis, we believe that there are still many open challenges and potentials for further advancements in this field. In this section, we list some future research directions for further exploration.
@@ 6.1 Efficiency
Notwithstanding the stellar progress and accomplishments of FMs, the improvements of performance in these models come at the expense of the model’s efficiency. However, anomaly detection models need to be particularly efficient because they are often applied in real-time, high-stakes environments where timely detection and response to anomalies are critical. Common application scenarios include fraud detection and healthcare, where even small delays of identification may lead to significant financial loss or even safety risks. While some lightweight adaptation techniques have been proposed to improve the parameter efficiency, Lester et al. (lester2021power, ) give an example that, as the model size increases, the performance gap between full fine-tuning and lightweight adaptation is diminishing rapidly. Therefore, exploring mechanisms to optimize the trade-off between efficiency and expressivity of FMs still remains a notable challenge.
@@ 6.2 Bias
In recent years, FMs have led to an extraordinary level of homogenization, which refers to the consolidation of methodologies for building machine learning systems across a wide range of applications (bommasani2021opportunities, ): current NLP models are often adapted from one of popular FMs like BERT. Despite that any improvements of FMs can help produce immediate benefits, such kind of extension might also have the potential to inherit or even amplify the problematic biases of these models. A biased anomaly detection model may potentially lead to incorrect or unfair outcomes. For example, a biased model may disproportionately flag specific data points or groups (e.g., gender and racial bias) as anomalous even when they are normal, limiting their capability of detecting genuine anomalies. Such kind of false positives or false negatives can lead to unfair treatment of certain individuals or groups, missed critical alerts, or a lack of trust in the system, ultimately compromising its effectiveness and fairness. Therefore, it is critical to eliminate the intrinsic bias present within FMs, thereby further avoiding extrinsic harms in downstream tasks.
@@ 6.3 Explainability
Despite that FMs are currently used as interpreters, in many works, to illustrate why anomalies are detected by providing textual explanations in a human-understandable manner, the models themselves are not transparent or interpretable enough to give explanations of the internal model structures and behaviors. However, providing evidence and logical steps for decision-making is a critical issue in anomaly detection especially in applications like healthcare and finance, where understanding the reasoning behind an anomaly is as important as detecting it (jin2025logicad, ). Therefore, how to improve the interpretability of FMs remains an open research question, and it is absolutely important for researchers to enhance the trustworthiness, usability, and decision-making transparency of anomaly detection.
@@ 6.4 Multimodality
While multimodality is considered a critical element of intelligence, and serves as a crucial component for the development of both thorough and broad comprehension of the world, models that go beyond simple alignment of vision and language are yet to emerge (radford2021learning, ; an2025gpt, ). In real-world application scenarios, taking healthcare as an example, medical data are highly multimodal, with various data types, scales, and styles. Multimodal models generally have better performance on anomaly detection by integrating and analyzing data from different sources than models with a single modality. However, current anomaly detection models are mainly developed for single modality (e.g., text, image, and gene), and do not learn from various modalities. By harnessing information across different data types, multimodal data can provide richer and more robust anomaly detection, thereby improving the accuracy of detection results across various applications. Future studies should further examine the design of FMs that integrate various modalities and domains.
@@ 7 Conclusion
The application of FMs to anomaly detection has emerged as a prominent area of research in recent years. In this survey, we provided an in-depth review of the use of FMs in anomaly detection tasks. We introduced a novel taxonomy classifying the methods into three categories based on the roles FMs may play in different stages of anomaly detection, namely encoder, detector, and interpreter. This taxonomy clarifies how FMs can empower the process of data representations, anomaly detection, and explainable analysis of detection results. We also discussed challenges and highlighted several future research directions, aiming to shed light on the way ahead in the field of anomaly detection with FMs. As FMs continue to evolve, their role in anomaly detection is expected to expand, unlocking new possibilities for more accurate, interpretable, and scalable anomaly detection systems.